# Wasserstein Spatial Depth

### v1

@@ Abstract
62G05,
60F15,
62G20,
Depths,
Measure Transportation,
@@ 1 Introduction
In this paper, we focus on on modeling distribution-valued data objects within Wasserstein spaces. There are several advantages to model certain data objects as distributions or probability measures. Firstly, it captures the hierarchical variations in the data by simulating a two-stage data-generating process: initially sampling multiple distributions from a Wasserstein space, followed by drawing data points from each sampled distribution Secondly, it captures variations of the data along geodesics of the distribution space that are not straight lines as in the Euclidean setting and thus are closer to the observations. Thirdly, it often provides a low-dimensional embedding that effectively represents high-dimensional data, enabling better statistical inference without the curse of the dimension. Since the Wasserstein space has different structure and property from the Euclidean space, conventional analytic tools cannot apply to distribution-valued data objects. Therefore, new methods specifically designed for analyzing such data are essential.
@@
There have been some efforts in this line of research, including but not limited to histogram regression [7], Wasserstein regression [12], geodesic PCA [5], template estimation [6], and Wasserstein clustering [48] or [18]. Despite the above development, there is limited effort in agnostic exploratory analysis for data objects in Wasserstein spaces [20, 16, 44, 19]. Still, exploratory analysis and descriptional statistics are critical to overview the properties of the data distribution before modeling. In particular, a notion of “ordering” for distribution-valued objects in Wasserstein spaces will be of fundamental utility. Besides exploratory analysis, it will also facilitate nonparametric methods for distribution-valued data.
@@
Quantiles, ranks, and signs are pivotal tools of semiparametric and nonparametric statistics. Due to the lack of canonical ordering in multi-dimensional Euclidean space, quantile or rank based tools have been limited to one-dimensional data before the creation of statistical depths. The notion of statistical depths fills this gap, extending the notion of order to higher dimension. Given a distribution $P$ on $\mathbb{R}^{d}$, the depth of any data point $\mathbf{x}\in\mathbb{R}^{d}$ is a non-negative value that measures the “centrality” of $\mathbf{x}$ with respect to $P$. A larger value of depth indicates the data point is more central within the distribution, while data points with small depths are considered outliers or less typical within the distribution, worthy of investigation. Several different types of depths have been proposed, including Tukey depth [39, 33], simplicial depth [25], spatial depth [11, 41], Monge-Kantorovich depth [13, 22] and lens depth [26]. Via endowing multivariate data points with “center-outward” orderings, depths allow extension of order statistics, robust inferences and classifications to multivariate data [49, 30].
@@
The concept of statistical depths has been extended to functional data, e.g., objects lying within a reproducing kernel Hilbert space (RKHS) [27, 32]. While it may be tempting to embed the Wasserstein space into a reproducing kernel Hilbert space (RKHS) and apply existing functional depth measures, this approach neglects the intricate geodesic structure of the Wasserstein space. There is no linear representation of the Wasserstein distance between distributions on $\mathbb{R}^{d}$ with $d>1$ [3]. Existing depths do not generalize well to nonlinear spaces. Besides the nonlinearity, the Wasserstein distance is computationally expensive even for empirical measures [35], which essentially rules out practical implementation of Tukey depth [39] and Monge-Kantorovich depth [13, 22]. The computational complexity of these two depths grows exponentially with the sample size.
@@ 1.1 Contributions
$\mathrm{SD}(\mathbf{x};P)=1-\left\|\mathbb{E}\left[\frac{\mathbf{X}-\mathbf{x}}{\|\mathbf{x}-\mathbf{X}\|}\right]\right\|,\quad\mathbf{X}\sim P.$
The spatial depth has been generalized to Hilbert spaces by following exactly the same definition [37, 41, 44]. However, the lack of linear structure of the Wasserstein space prevents an straightforward adaptation of the spatial depth. Nevertheless, the Wasserstein metric endows the space of probability measures with a structure of geodesic metric space (see [1]). For absolutely continuous probability measures $Q$ and $P$, the constant speed geodesic joining $Q$ and $P$ is given by the curve of probability measures
@@
Above, $T_{Q,P}$ is the optimal transport map from $Q$ to $P$ and $\mathcal{W}_{2}(P,Q)$ is the Wasserstein distance between $P$ and $Q$, both being formally defined in Section 3.3. We will show that $\mathrm{SD}(Q;\mathbf{P})$ satisfies the same properties as its Euclidean counterpart, namely transformation invariance, taking values in $[0,1]$, decreasing at infinity and attaining maximum at the median. Moreover, we show that $Q\mapsto\mathrm{SD}(Q;{\bf P})$ is continuous and ${\bf P}\mapsto\mathrm{SD}(Q;{\bf P})$ is continuous. Next, we will propose a finite-sample estimator under the so-called two-stage and one-stage sampling models. In the two-stage sampling model, such an estimator can be computed in polynomial time. Moreover, we will show that in both models, the estimator is consistent, meaning that it approximates the true population depth function as the sample size increases. We also prove asymptotic normality. Finally, we will provide numerical simulations for real and synthetic datasets.
@@ 1.2 Organization
General notations are provided in Section 2. The definition of WSD is given in Section 3 with illustrating examples in Section 4. In Section 5, we show that WSD shares the desirable properties of conventional statistical depths [49]. In Section 6, we tackle consistent estimation with asymptotic normality. In Section 7, we compare WSD to several depths in general metric spaces [20, 16, 44] adapted to Wasserstein spaces. We advocate WSD over the other depths in terms of computational feasibility, assumption flexibility, while possessing all desirable properties of a depth. In section 8, extensive numerical simulations are shown to demonstrate the empirical validity of WSD. Finally, in Section 9, we apply it to explore real-world data and make informative discoveries. All the proofs are provided in the appendix.
@@ 2 Notation
denotes the $L^{2}(P)$-norm of $f$. The Hilbert space of measurable functions with finite $L^{2}(P)$-norm is denoted as $L^{2}(P)$ with inner product $\langle\cdot,\cdot\rangle_{L^{2}(P)}$. We say that a sequence $\{\mu_{n}\}_{n\in\mathbb{N}}\subset\mathcal{P}(\mathcal{K})$ converges weakly to $\mu\in\mathcal{P}(\mathbb{X})$ if
@@
$d_{{\rm BL}}(\mu,\nu)\\ =\sup\bigg{\{}\int f(x)\mathrm{d}(\mu-\nu)(x):\ |f(x)|\leq 1\ {\rm and}\ |f(x)-f(y)|\leq d(x,y),\ \forall\,x,y\in\mathbb{X}\bigg{\}}.$
@@ 3 From Euclidean to Wasserstein spatial depth
In this section we define our notion of Wasserstein spatial depth. In Section 3.1 we recall the definition of Euclidean spatial depth and its main properties. In Section 3.2 we provide our interpretation of spatial depth in terms of geodesics, which allows its generalization to the Wasserstein space of measures (see Section 3.3). For readers interested in a more comprehensive understanding of the mathematical concepts discussed in Sections 3.2 and 3.3, we recommend consulting the monograph [1] for an in-depth exposition.
@@ 3.1 Euclidean spatial depth
Throughout the paper, we use the convention $/0=0$ 0. The spatial depth shares the following properties with the univariate canonical depth function $2\min(F(x),1-F(-x))$. First, the statistical depth $\mathrm{SD}(\mathbf{x};P)$ belongs to the interval $[0,1]$. Second, the geometric median, defined as
@@
$\mathrm{SD}(T(\mathbf{x});T\#P)=\mathrm{SD}(\mathbf{x};P).$
@@ 3.3 Geodesic spatial depth over the space of measures
$\mathcal{T}_{p}(P,Q)=\inf_{\pi\in\Pi(P,Q)}\frac{1}{2}\int\|\mathbf{x}-\mathbf{y}\|^{p}\mathrm{d}\pi(\mathbf{x},\mathbf{y}),$ (1)
where $\Pi(P,Q)\subset\mathcal{P}_{p}(\mathbb{R}^{d}\times\mathbb{R}^{d})$ stands for the set of probability measures with marginals $P$ and $Q$, i.e., $(\mathbf{X},\mathbf{Y})\sim\pi\in\Pi(P,Q)$ if $\mathbf{X}\sim P$ and $\mathbf{Y}\sim Q$. For $p\geq 1$, the mapping $(P,Q)\mapsto\mathcal{W}_{p}(P,Q)=\left(\mathcal{T}_{p}(P,Q)\right)^{\frac{1}{p}}$ defines a distance over the space $\mathcal{P}_{p}(\mathbb{R}^{d})$ such that
@@
We focus now on the case $p=2$. We define $\mathcal{P}^{a.c}_{2}(\mathbb{R}^{d})$ as the subset of $\mathcal{P}_{2}(\mathbb{R}^{d})$ composed of absolutely continuous measures. If $P$ belongs to $\mathcal{P}^{a.c}_{2}(\mathbb{R}^{d})$, there exists a unique minimizer $\pi_{P,Q}$ of (1), for $p=2$. Moreover, there exists a unique gradient of a convex function $T_{P,Q}=\nabla\phi_{P,Q}$ such that $\pi_{P,Q}=({\rm I}\times\nabla\phi_{P,Q})_{\#}P$. The map $T_{P,Q}$ is called an optimal transport map. Here, for a probability measure $\mu$ and a Borel mapping $T$, $T_{\#}\mu$ denotes the push forward measure, which is the distribution of $T(\mathbf{X})$, for $\mathbf{X}\sim\mu$.
@@
$\partial_{t}P_{t}+{\rm div}({\bf v}_{t}P_{t})=0$ (2)
in an appropriate weak sense. The continuity equation is commonly used in fluid mechanics, where ${\bf v}_{t}$ represents the flow velocity vector field. However, given the curve $\{\mathbf{X}_{t}\}_{t\in[0,1]}$ there could exist several curves of velocity fields $\{{\bf v}_{t}\}_{t\in[0,1]}$ solving (2), i.e., generating the same flow. Among all of them, there exists only one belonging to
$\operatorname*{arg\,min}\left\{\int_{0}^{1}\|{\bf v}_{t}\|_{L^{2}(P_{t})}^{2}\mathrm{d}t:\quad\partial_{t}P_{t}+{\rm div}({\bf v}_{t}P_{t})=0\right\}.$ (3)
@@
$\operatorname*{arg\,min}\left\{\int_{0}^{1}\|{\bf v}_{t}\|_{L^{2}(\gamma_{t})}^{2}\mathrm{d}t:\quad\partial_{t}\gamma_{t}+{\rm div}({\bf v}_{t}\gamma_{t})=0,\ \gamma_{0}=P\ {\rm and}\ \gamma_{1}=Q\right\}.$ (4)
@@
Since, ${\bf v}_{0}^{Q\to P}=T_{Q\to P}-{\rm I}$ we get the following definition of spatial depth.
@@ 4.1 Univariate case
In the case of univariate distributions, WSD reduces to quantile spatial depth. The univariate Wasserstein distance has a flat structure since there is an isometric homeomorphism between distributions and the corresponding CDF’s. Consequently, the Wasserstein distance between univariate distributions $P$ and $Q$ has a simple form
@@
$\mathrm{SD}(Q;{\bf P}):=1-\left(\int_{0}^{1}{\left(\mathbb{E}_{P\sim\mathbf{P}}\left[\frac{F^{-1}_{P}(u)-F^{-1}_{Q}(u)}{\left(\int_{0}^{1}(F^{-1}_{P}(u)-F^{-1}_{Q}(u))^{2}\mathrm{d}\ell_{1}(u)\right)^{\frac{1}{2}}}\right]\right)^{2}}\mathrm{d}\ell_{1}(u)\right)^{\frac{1}{2}},$
@@ 4.2 Location families
When $\mathbf{P}$ is supported on a location family, a set of shifted distributions indexed by the location parameter $\boldsymbol{\theta}$. In this case, $\mathbf{P}$ coincides with the distribution of the location parameter. And the WSD reduces to
$\mathrm{SD}(Q;{\bf P})=1-\left\|\mathbb{E}_{P\sim\mathbf{P}}\Bigg{(}\frac{\boldsymbol{\theta}_{P}-\boldsymbol{\theta}_{Q}}{\|\boldsymbol{\theta}_{P}-\boldsymbol{\theta}_{Q}\|}\Bigg{)}\right\|=1-\left\|\mathbb{E}_{\boldsymbol{\theta}}\Bigg{(}\frac{\boldsymbol{\theta}-\boldsymbol{\theta}_{Q}}{\|\boldsymbol{\theta}-\boldsymbol{\theta}_{Q}\|}\Bigg{)}\right\|,$ (5)

### v2

@@ Abstract
62R10,
62G30,
62G35,
Distributional data analysis,
High dimensional data,
Order statistic,
@@ 1 Introduction
In this paper, we focus on modeling distribution-valued data objects within Wasserstein spaces. There are several advantages to model certain data objects as distributions or probability measures. Firstly, it captures the hierarchical variations in the data by simulating a two-stage data-generating process: initially sampling multiple distributions from a Wasserstein space, followed by drawing data points from each sampled distribution. Secondly, it captures variations of the data along geodesics of the distribution space that are not straight lines as in the Euclidean setting and thus are closer to the observations. Thirdly, it often provides a low-dimensional embedding that effectively represents high-dimensional data, enabling better statistical inference without the curse of the dimension. Since the Wasserstein space has different structure and property from the Euclidean space, conventional analytic tools cannot apply to distribution-valued data objects. Therefore, new methods specifically designed for analyzing such data are essential.
@@
There have been some efforts in this line of research, including but not limited to histogram regression [9], Wasserstein regression [2, 15], geodesic PCA [7], template estimation [8], and Wasserstein clustering [62, 24]. Despite the above developments, there is limited effort in agnostic exploratory analysis for data objects in Wasserstein spaces [28, 22, 59, 25]. Still, exploratory analysis and descriptional statistics are critical to overview the properties of the data distribution before modeling. In particular, a notion of “ordering” for distribution-valued objects in Wasserstein spaces will be of fundamental utility. Besides exploratory analysis, it will also facilitate nonparametric methods for distribution-valued data.
@@
Quantiles, ranks, and signs are pivotal tools of semiparametric and nonparametric statistics. Due to the lack of canonical ordering in multi-dimensional Euclidean space, quantile or rank based tools have been limited to one-dimensional data before the creation of statistical depths. The notion of statistical depths fills this gap, extending the notion of order to higher dimension. Given a distribution $P$ on $\mathbb{R}^{d}$, the depth of any data point $\mathbf{x}\in\mathbb{R}^{d}$ is a non-negative value that measures the “centrality” of $\mathbf{x}$ with respect to $P$. A larger value of depth indicates the data point is more central within the distribution, while data points with small depths are considered outliers or less typical within the distribution, worthy of investigation. Several different types of depths have been proposed, including Tukey depth [54, 47], simplicial depth [34], spatial depth [14, 56], Monge-Kantorovich depth [16, 31] and lens depth [35]. Via endowing multivariate data points with “center-outward” orderings, depths allow extension of order statistics, robust inference [36, 63] and classification to multivariate data [64, 43].
Statistical depth theory is one of the main research areas of functional data analysis (FDA). Most of the Euclidean depth functions extend naturally to Hilbert-space-valued data, see [38, 46, 45, 20]. For instance, this is the case for the h-depth [19], the Tukey depth [26] and its random version [18], the spatial depth [11], the integrated depth [20] and the Monge-Kantorovich depth [30]. For Banach spaces, some examples are the integrated depth [20], the band depth and its modified version [38], the half-region depth and its modified version [39], the $L^{\infty}$ depth [37] and the infimal depth [42].
@@
While it may be tempting to embed the Wasserstein space into a function space, for instance a reproducing kernel Hilbert space (RKHS) [52], and apply existing functional depth measures, this approach neglects the intricate geodesic structure of the Wasserstein space. There is no linear representation of the Wasserstein distance between distributions on $\mathbb{R}^{d}$ with $d>1$ [5]. Existing depths do not generalize well to nonlinear spaces. Besides the nonlinearity, the Wasserstein distance is computationally expensive even for empirical measures [49], which essentially rules out practical implementation of Tukey depth [54] and Monge-Kantorovich depth [16, 31]. The computational complexity of these two depths grows exponentially with the sample size.
@@ 1.1 Contributions
$\mathrm{SD}(\mathbf{x};P)=1-\left\|\mathbb{E}\left[\frac{\mathbf{X}-\mathbf{x}}{\|\mathbf{X}-\mathbf{x}\|}\right]\right\|,\quad\mathbf{X}\sim P.$ (1.1)
The spatial depth has been generalized to Hilbert spaces by following exactly the same definition [51, 56, 59]. However, the lack of linear structure of the Wasserstein space prevents a straightforward adaptation of the spatial depth. Nevertheless, the Wasserstein metric endows the space of probability measures with a structure of geodesic metric space (see [1]). For absolutely continuous probability measures $Q$ and $P$, the constant speed geodesic joining $Q$ and $P$ is given by the curve of probability measures
@@
Above, $T_{Q,P}$ is the optimal transport map from $Q$ to $P$ and $\mathcal{W}_{2}(P,Q)$ is the Wasserstein distance between $P$ and $Q$, both being formally defined in Section 3.3. We will show that $\mathrm{SD}(Q;\mathbf{P})$ satisfies the same properties as its Euclidean counterpart, namely transformation invariance, taking values in $[0,1]$, decreasing at infinity and attaining maximum at the median. Moreover, we show that $Q\mapsto\mathrm{SD}(Q;{\bf P})$ is continuous and ${\bf P}\mapsto\mathrm{SD}(Q;{\bf P})$ is continuous. Next, we will propose a finite-sample estimator under the so-called two-stage and one-stage sampling models. In the two-stage sampling model, such an estimator can be computed in polynomial time. Moreover, we will show that in both models, the estimator is consistent, meaning that it approximates the true population depth function as the sample size increases. We also prove asymptotic normality. Finally, we will provide numerical simulations for real and synthetic datasets. In particular, we highlight that our suggested depth is more informative than depth methods designed for linear spaces and applied to mappings of distributions to these linear spaces.
@@ 1.2 Organization
General notations are provided in Section 2. The definition of WSD is given in Section 3 with illustrating examples in Section 4. In Section 5, we show that WSD shares the desirable properties of conventional statistical depths [64]. In Section 6, we tackle consistent estimation with asymptotic normality. In Section 7, we compare WSD to several depths in general metric spaces [28, 22, 59] adapted to Wasserstein spaces. We advocate WSD over the other depths in terms of computational feasibility and assumption flexibility, while possessing all desirable properties of a depth. In section 8, extensive numerical simulations are shown to demonstrate the empirical validity and merits of WSD. Finally, in Section 9, we apply it to explore real-world data and make informative discoveries. All the proofs are provided in the Appendix.
@@ 2 Notation
denotes the $L^{2}(P)$-norm of $f$. The Hilbert space of measurable functions with finite $L^{2}(P)$-norm is denoted as $L^{2}(P)$ with inner product $\langle\cdot,\cdot\rangle_{L^{2}(P)}$. We also extend the definition of the Hilbert space $L^{2}(P)$ and the associated notation to vector-valued functions, with for $f,g:\mathcal{K}\to\mathbb{R}^{k}$,
$\langle f,g\rangle_{L^{2}(P)}:=\int\langle f,g\rangle\mathrm{d}P.$
We say that a sequence $\{\mu_{n}\}_{n\in\mathbb{N}}\subset\mathcal{P}(\mathcal{K})$ converges weakly to $\mu\in\mathcal{P}(\mathcal{K})$ if
@@
$d_{{\rm BL}}(\mu,\nu)=\sup\bigg{\{}\int f(x)\mathrm{d}(\mu-\nu)(x):\\ |f(x)|\leq 1\ {\rm and}\ |f(x)-f(y)|\leq d(x,y),\ \forall\,x,y\in\mathcal{K}\bigg{\}}.$
@@ 3 From Euclidean to Wasserstein spatial depth
In this section we define our notion of Wasserstein spatial depth. In Section 3.1 we recall the definition of Euclidean spatial depth and its main properties. In Section 3.2 we provide our interpretation of spatial depth in terms of geodesics, which allows for its generalization to the Wasserstein space of measures (see Section 3.3). For readers interested in a more comprehensive understanding of the mathematical concepts discussed in Sections 3.2 and 3.3, we recommend consulting the monograph [1] for an in-depth exposition.
@@ 3.1 Euclidean spatial depth
[…]
@@
$\mathrm{SD}(T(\mathbf{x});T\#P)=\mathrm{SD}(\mathbf{x};P),$
where again the push-forward operator ##\## is defined in Section 3.3.
@@ 3.3 Geodesic spatial depth over the space of measures
${\rm OT}_{p}(P,Q)=\inf_{\pi\in\Pi(P,Q)}\frac{1}{2}\int\|\mathbf{x}-\mathbf{y}\|^{p}\mathrm{d}\pi(\mathbf{x},\mathbf{y}),$ (3.1)
where $\Pi(P,Q)\subset\mathcal{P}_{p}(\mathbb{R}^{d}\times\mathbb{R}^{d})$ stands for the set of probability measures with marginals $P$ and $Q$, i.e., $(\mathbf{X},\mathbf{Y})\sim\pi\in\Pi(P,Q)$ if $\mathbf{X}\sim P$ and $\mathbf{Y}\sim Q$. For $p\geq 1$, the mapping $(P,Q)\mapsto\mathcal{W}_{p}(P,Q)=\left({\rm OT}_{p}(P,Q)\right)^{\frac{1}{p}}$ defines a distance over the space $\mathcal{P}_{p}(\mathbb{R}^{d})$ such that
@@
We focus now on the case $p=2$. We define $\mathcal{P}^{a.c}_{2}(\mathbb{R}^{d})$ as the subset of $\mathcal{P}_{2}(\mathbb{R}^{d})$ composed of absolutely continuous measures. If $P$ belongs to $\mathcal{P}^{a.c}_{2}(\mathbb{R}^{d})$, there exists a unique minimizer $\pi_{P,Q}$ of (3.1), for $p=2$. Moreover, there exists a unique gradient of a convex function $T_{P,Q}=\nabla\phi_{P,Q}$ such that $\pi_{P,Q}=({\rm I}\times\nabla\phi_{P,Q})_{\#}P$. The map $T_{P,Q}$ is called an optimal transport map. Here, for a probability measure $\mu$ and a Borel mapping $T$, $T_{\#}\mu$ denotes the push forward measure, which is the distribution of $T(\mathbf{X})$, for $\mathbf{X}\sim\mu$.
@@
$\partial_{t}P_{t}+{\rm div}({\bf v}_{t}P_{t})=0$ (3.2)
in an appropriate weak sense. The continuity equation is commonly used in fluid mechanics, where ${\bf v}_{t}$ represents the flow velocity vector field. However, given the curve $\{\mathbf{X}_{t}\}_{t\in[0,1]}$ there could exist several curves of velocity fields $\{{\bf v}_{t}\}_{t\in[0,1]}$ solving (3.2), i.e., generating the same flow. Among all of them, there exists only one belonging to
$\operatorname*{arg\,min}\left\{\int_{0}^{1}\|{\bf v}_{t}\|_{L^{2}(P_{t})}^{2}\mathrm{d}t:\quad\partial_{t}P_{t}+{\rm div}({\bf v}_{t}P_{t})=0\right\}.$ (3.3)
@@
$\operatorname*{arg\,min}\left\{\int_{0}^{1}\|{\bf v}_{t}\|_{L^{2}(\gamma_{t})}^{2}\mathrm{d}t:\quad\partial_{t}\gamma_{t}+{\rm div}({\bf v}_{t}\gamma_{t})=0,\ \gamma_{0}=P\ {\rm and}\ \gamma_{1}=Q\right\}.$ (3.4)
@@
Since ${\bf v}_{0}^{Q\to P}=T_{Q,P}-{\rm I}$ we get the following definition of spatial depth.
@@ 4.1 Univariate case
In the case of univariate distributions, WSD reduces to quantile spatial depth. The univariate Wasserstein distance has a flat structure since there is an isometric homeomorphism between distributions and the corresponding generalized quantile functions. Consequently, the Wasserstein distance between univariate distributions $P$ and $Q$ has a simple form
@@
$\mathrm{SD}(Q;{\bf P}):=1-\left(\int_{0}^{1}{\left(\mathbb{E}_{P\sim\mathbf{P}}\left[\frac{F^{-1}_{P}(u)-F^{-1}_{Q}(u)}{\left(\int_{0}^{1}(F^{-1}_{P}(u)-F^{-1}_{Q}(u))^{2}\mathrm{d}u\right)^{\frac{1}{2}}}\right]\right)^{2}}\mathrm{d}u\right)^{\frac{1}{2}},$
@@ 4.2 Location families
Consider that $\mathbf{P}$ is supported on a location family, a set of shifted distributions indexed by the location parameter $\boldsymbol{\theta}$. In this case, $\mathbf{P}$ coincides with the distribution of the location parameter. And the WSD reduces to
$\mathrm{SD}(Q;{\bf P})=1-\left\|\mathbb{E}_{P\sim\mathbf{P}}\Bigg{(}\frac{\boldsymbol{\theta}_{P}-\boldsymbol{\theta}_{Q}}{\|\boldsymbol{\theta}_{P}-\boldsymbol{\theta}_{Q}\|}\Bigg{)}\right\|=1-\left\|\mathbb{E}_{\boldsymbol{\theta}}\Bigg{(}\frac{\boldsymbol{\theta}-\boldsymbol{\theta}_{Q}}{\|\boldsymbol{\theta}-\boldsymbol{\theta}_{Q}\|}\Bigg{)}\right\|,$ (4.1)